----------------------------------------------------------------------------- Mauna Loa Solar Observatory Observer's Log ----------------------------------------------------------------------------- Tue Oct 20 16:38:50 GMT 2026 Year:26 Doy:293 Observer: mcotter Weather COMMENT BY mcotter: Tue Oct 20 16:44:15 GMT 2026 MLSO Current Weather Timestamp (UTC): 2026-10-20 16:35:00 Temperature : 42.12 deg F Dew Point : 2.61 deg F Pressure : 681.44 mbar Wind Speed : 9.4 mph Wind Direction : 143.7 deg Skies: Clear skies overhead But very large bands of Cirrocumulus clouds are present on the north, west and eastern horizons. The cloud bands to the east look like they trail off to the south. Very hazy conditions at lower elevations with the inversion layer visible on the horizon just at the peak of Hualalai. Light steady wind blowing from the southeast. ___end___ Source Log Type BY mcotter: Tue Oct 20 16:48:32 GMT 2026 entry was edited noaa construction: The NOAA construction contractors are now onsite and work has already started for the day with heavy equipment starting to traverse the area. I was told by the crew supervisor that todays activities will involve setting up cement forms with the likelihood of heavy equipment being used by the crew throughout the morning into the early afternoon. The wind is currently blowing from the south so we should be able to observe as long as conditions allow. ___end___ GENERAL COMMENT BY mcotter: Tue Oct 20 17:00:53 GMT 2026 Opened windows upstairs ___end___ GENERAL COMMENT BY mcotter: Tue Oct 20 17:00:58 GMT 2026 Dome open ___end___ ** KCOR PM mcotter ** : Tue Oct 20 17:01:05 GMT 2026 Blew off Kcor O1 ___end___ ** UCOMP PM mcotter ** : Tue Oct 20 17:01:12 GMT 2026 Blew off UCoMP O1 ___end___ BY mcotter: Tue Oct 20 17:14:55 GMT 2026 Weather Report: Weather satellite images show very large bands of upper level clouds moving towards the Big Island from the southwest. It appears that it will arrive later this morning or in the early afternoon. ___end___ Tue Oct 20 17:20:23 GMT 2026 Kcor Focus/alignment program exited Tue Oct 20 17:23:45 GMT 2026 Kcor Focus/alignment program exited ** KCOR PROBLEM REPORTED BY mcotter ** : Tue Oct 20 19:31:55 GMT 2026 This morning while observing the bottom two images in the synoptic GUI turned completely white. The top eight synoptic images looked slightly bright but were completely acceptable. The NRGF and Level-0 images appeared strange, making me suspect a circular acquisition error. I stopped the observing code and the socketcam GUI, but the command window showed no indication of an acquisition error. I restarted the observing program and socketcam and again the bottom two synoptic images were all white, but again the top eight synoptic images looked acceptable. When the data appeared in the NRGF and Level-0 images they again looked like there was an acquisition error. Ben checked to ensure everything was operating correctly (and it was) but determined that today's sky conditions were causing polarization issues with the data. Today, high-level Cirrus and Cirrostratus clouds are spread across the sky in different areas. Although the area we were observing looked clear, a very thin translucent layer of clouds must still be present. This thin layer of clouds combined with the low angle of our observations of the Sun appears to have caused an unwanted polarization in the data even though our initial synoptic images looked good. Due to poor sky conditions observations have been stopped. ___end___ Weather COMMENT BY mcotter: Tue Oct 20 19:59:35 GMT 2026 MLSO Current Weather Timestamp (UTC): 2026-10-20 19:53:00 Temperature : 55.27 deg F Dew Point : -8.28 deg F Pressure : 682.31 mbar Wind Speed : 11.41 mph Wind Direction : 66.1 deg Skies: The sky is now completely overcast with Cirrus and Cirrostratus clouds. A moderate steady wind continues to blow from the east-northeast. ___end___ BY mcotter: Tue Oct 20 20:00:30 GMT 2026 Due to poor sky conditions the dome area windows and the dome shutters are now closed. ___end___ ** KCOR PM mcotter ** : Wed Oct 21 02:32:53 GMT 2026 Kcor: First Contact was applied to the opposite side of the #4-O1 Lens this afternoon; yesterday we had applied First Contact to the first side of the Kcor #4-O1 Lens. We are going to allow the First Contact coatings to fully cure before removing the applications. ___end___ BY mcotter: Wed Oct 21 02:37:06 GMT 2026 End Of Day: Some observations were accomplished in the early morning but within a short period of time large bands of high level clouds came into the observing area and observations were stopped. Weather satellite images for the next few days do not look promising for taking observations, but there may be breaks in the cloud coverage that provide viewing opportunities. ___end___ Logger COMMENT BY mcotter: Wed Oct 21 02:37:11 GMT 2026 Log closed ONSITE STAFF: berkey, mcotter ___end___ KCOR COMMENT BY KCOR: Wed Oct 21 02:37:52 GMT 2026 KCOR file transfer report KCor created 245 .bin files 245 .fts files Remote site is complete ? no rsync required. ___end___ Logger COMMENT BY : Wed Oct 21 15:39:10 GMT 2026 Log closed ONSITE STAFF: ___end___